Grade 11Mechanics

A body starts from rest with an acceleration of 2 metre per second square till it attains the maximum velocity then retards to rest with 3 metre per second square if total time taken is 10 second maximum speed attains is

rubiks cuber
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student

Here,

u = 0

a1 = 2 ms-2

t = 10 s

a2 = -3ms-2

Let the maximum velocity attained be v at time t

Thus

v = u + a1t

=> v = 2t  ---1.

Again it reaches velocity 0 after 10 – t seconds with a retardation of 3 ms-2

Thus

0 = v +a2×(10-t)

=> 0 = v – 3×(10-t)

By 1 we have

=> 2t – 3(10-t) = 0

=> 5t = 30

=> t = 6 s

Again,

v = 2t

= 2×6 = 12 ms-1  

A body starts from rest with an acceleration 2m/s2 till it attains the maximum velocity then retards to rest with 3m/s2. If total time taken is 10 seconds , then maximum speed attained is____12_m/s__ .


Regards

Arun (askIITians forum expert)